OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 5t − 7(3 + t) 12t − 21 12t − 14 −2t − 21 −2t + 21

OpenStudy (lgbasallote):

-2t -21

OpenStudy (anonymous):

yeaup!

Directrix (directrix):

5t − 7(3 + t) = 5t - (7)(3) - (7)(t) = 5t - 21 - 7t = -2 t - 21